# Bayesian Model Averaging in Causal Instrumental Variable Models

### v4

@@ 3.3 Model Space
$p(L_{j})=\frac{\Gamma(a+b_{L})}{\Gamma(a)\Gamma(b_{L})}\frac{\Gamma(a+k_{j})\Gamma(b_{L}+k-k_{j})}{\Gamma(a+b_{L}+k)}.$
Following Ley and Steel, (2009), we choose $a=1$ and $b_{L}=(k-m_{L})/m_{L}$, where the user specifies the prior mean outcome model size $m_{L}$. If not stated otherwise, we use $m_{L}=p/2$ as our default choice. For the treatment model we use a similar procedure, leading to uniform hyperpriors on the prior inclusion probabilities, as in Scott and Berger, (2010).
@@ 4.1 Computational Strategy
    Sample the covariance matrix $\bm{\Sigma}^{(s)}$ from its full conditional distribution in (7).
@@
where $h$ is a proposal kernel. Throughout, models are proposed by randomly permuting the inclusion index of one covariate. This proposal is symmetric in the sense that $h(L\mid L^{\prime})/h(L^{\prime}\mid L)=1$, so the acceptance probability reduces to the (conditional) Bayes factor times the prior ratio (capped at 1). This can lead to slow mixing in very high dimensions, where other proposals might be more attractive, but we leave this to future work.
@@ 4.2 Conditional Bayes factors and posteriors
Finally, we have the conditional posterior for the covariance matrix,
$\bm{\Sigma}\mid\bm{\rho,\Lambda},L,M,\bm{y,X}\sim W^{-1}(\nu+n,\bm{I_{l+1}}+\bm{[\epsilon:H]}^{\intercal}\bm{[\epsilon:H]}).$ (7)
@@ 5 Model Selection consistency of gIVBMA
  • under a hyperprior $p(g_{M})$, we have for $c\in\{l,2l,\dots,(p+k)l\}$
@@ 6 Simulation experiments
We evaluate the performance of gIVBMA in three different scenarios and compare it to naive BMA, IVBMA, and several classical methods. Throughout, we consider the median absolute error (MAE) and median bias of the point estimates of $\bm{\tau}$, (credible or confidence) interval coverage of $\bm{\tau}$, and the log predictive score (LPS) on a separate holdout dataset. For all Bayesian methods, we use the posterior mean as our point estimate for the MAE and bias calculations. While MAE, bias, and coverage relate specifically to the quality of the estimation of $\bm{\tau}$, LPS is a measure of predictive adequacy, focused on probabilistic predictions of the outcomes. Supplementary Section S.4.1 provides more details on the different measures used, while Section S.4.2 describes the different estimators and their implementation.
The predictive comparison is perhaps not entirely fair to some of the classical methods (in particular those that are two-stage in nature) as they reduce the bias in $\bm{\tau}$ but lack a correction term like gIVBMA. Consequently, when making predictions, these methods use an estimate of $\bm{\tau}$ instead of $\bm{\tau+\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$, which would be the appropriate coefficient for the conditional model $\bm{y}\mid\bm{X}$. This can lead to poor predictions when $\bm{\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$ becomes large. However, in practice, these methods might not be employed when prediction is the primary objective. Methods that do not correct for endogeneity (OLS or naive BMA) target the biased coefficient $\bm{\tau+\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$, but provide good predictions as this is the right coefficient for predicting $\bm{y}$ conditional on $\bm{X}$. This discrepancy reflects the fact that predicting observed and counterfactual outcomes are different problems under endogeneity. Our probabilistic method allows us to provide endogeneity-corrected inference on $\bm{\tau}$ while maintaining good predictive power in the observational model.
@@ 6.1 Invalid instruments
We compare the performance of our method to IVBMA (Karl and Lenkoski,, 2012), TSLS using all instruments, oracle TSLS, and the sisVIVE estimator of Kang et al., (2016). We also add what we call oracle gIVBMA (O-gIVBMA), which correctly separates the valid and invalid instruments (or covariates) a priori. For this oracle, the invalid instruments can be included in both models, but the valid instruments can only be included in the treatment model. This corresponds to a situation where the analyst has some extra prior information to distinguish valid instruments from potentially invalid instruments. We would expect this to be slightly more efficient in smaller samples. For this oracle version, we also consider a two-component prior (2C) as proposed in Zhang et al., (2016), and described in detail in supplementary Section S.1.
@@
BMA (hyper-$g/n$) 0.48 0.48 0.09 1.38 0.48 0.48 0.09 1.41
gIVBMA (BRIC) 0.35 0.34 1.0 1.37 0.26 0.25 0.98 1.41
gIVBMA (hyper-$g/n$) 0.13 0.09 0.97 1.37 0.2 0.0 0.97 1.41
O-gIVBMA (hyper-$g/n$) 0.16 0.09 0.99 1.37 0.14 -0.0 0.98 1.39
O-gIVBMA (2C) 0.14 0.07 1.0 1.37 0.17 -0.02 0.97 1.39
IVBMA (KL) 0.39 0.39 0.98 1.38 0.6 0.6 0.89 1.41
OLS 0.53 0.53 0.08 1.44 0.51 0.51 0.05 1.5
TSLS 1.45 1.45 0.21 2.15 2.44 2.44 0.03 2.46
O-TSLS 0.28 0.25 0.77 1.45 0.31 0.24 0.88 1.5
sisVIVE 0.66 0.66 - 1.69 2.41 2.41 - 2.39
@@
BMA (hyper-$g/n$) 0.46 0.46 0.0 1.29 0.48 0.48 0.0 1.27
gIVBMA (BRIC) 0.08 0.06 0.96 1.29 0.19 0.19 0.94 1.27
gIVBMA (hyper-$g/n$) 0.1 0.06 0.95 1.29 0.25 0.24 0.91 1.27
O-gIVBMA (hyper-$g/n$) 0.06 0.02 0.96 1.29 0.08 0.03 0.92 1.27
O-gIVBMA (2C) 0.07 0.03 0.96 1.29 0.1 0.04 0.93 1.27
IVBMA (KL) 0.09 -0.06 0.82 1.29 0.08 -0.05 0.83 1.27
OLS 0.5 0.5 0.0 1.29 0.5 0.5 0.0 1.27
TSLS 1.79 1.79 0.0 2.18 3.57 3.57 0.0 2.72
O-TSLS 0.06 0.02 0.96 1.42 0.08 0.02 0.95 1.41
sisVIVE 0.22 0.22 - 1.61 4.36 4.36 - 3.05
Table 1: Invalid Instruments: MAE, bias, coverage, and mean LPS on 100 simulated datasets with $s$ invalid instruments. The best values in each column are printed in bold. The sisVIVE estimator does not provide any uncertainty quantification, so we do not report any coverage results.
Table 1 shows that all gIVBMA variants predict well and lead to small bias and MAE. For small $n$, the BRIC specification has a higher MAE and median bias than the hyper-$g/n$ variants. The oracle versions tend to perform slightly better, suggesting that using prior information on the separation of instruments and covariates can be beneficial. The IVBMA method also predicts well, but incurs significant bias for small $n$ and undercovers for large $n$. TSLS results in a very large positive bias, very small coverage, and poor prediction. The oracle version of TSLS does better but is not competitive with the gIVBMA methods for small $n$. All the methods above appear largely unaffected by the plurality rule. The latter does affect sisVIVE, however, which performs very poorly when $s=6$, as expected.
@@ 6.2 Multiple endogenous variables with correlated instruments
We compare gIVBMA against naive BMA, IVBMA, OLS, TSLS (using all instruments and an oracle version), and MATSLS. The latter refers to the model-averaged TSLS estimator with unrestricted weights by Kuersteiner and Okui, (2010) and is included as it is the only classical method that is competitive with oracle TSLS in all scenarios of Subsection 6.3. To evaluate the model selection performance of gIVBMA and IVBMA, we also compare the distribution of the estimated posterior probabilities of the true treatment model $M=\{1,5,7,11,13\}$ and the mean treatment model sizes. IVBMA estimates separate treatment models for the two endogenous variables, and we report both. IVBMA cannot directly model the Beta-distributed endogenous variable and instead relies on a Gaussian approximation.
Both gIVBMA variants outperform all other methods in terms of MAE and median bias, even in the larger sample. With $n=500$, they slightly undercover for the Beta endogenous variable, while their coverage is too large for $n=50$. BMA and OLS are severely biased as expected, but their coverage for the second endogenous variable is adequate (in contrast with the coverage for $\bm{X}_{1}$). This is a consequence of the non-linear transformation from the latent Gaussian to the Beta variable, diluting the endogeneity. TSLS, oracle TSLS, and MATSLS all perform similarly well, with only some under-coverage for $n=50$. IVBMA results in a substantial bias for both sample sizes, yet it predicts and covers relatively well. Supplementary Table S.3 presents the full estimation error, coverage, and LPS results.
@@
Figure 1 presents the distributions of the posterior probability of the true treatment model and the mean treatment model size. Both gIVBMA variants put substantial posterior mass on the true model, and their mean model size is accordingly very close to that of the true model ($5$). For $n=50$, IVBMA tends to select too many variables for $X_{1}$ and too few for $X_{2}$. The performance of IVBMA is better in the $n=500$ setting, but it still cannot compete with gIVBMA. Supplementary Table S.4 presents median posterior inclusion probabilities (PIP) for all instruments. The selection performance of gIVBMA is clearly superior for both sample sizes.

### v5

@@ 3.3 The prior on the model space
$p(L_{j})=\frac{\Gamma(a+b_{L})}{\Gamma(a)\Gamma(b_{L})}\frac{\Gamma(a+p_{j})\Gamma(b_{L}+p-p_{j})}{\Gamma(a+b_{L}+p)}.$
Following Ley and Steel, (2009), we choose $a=1$ and $b_{L}=(p-m_{L})/m_{L}$, where the user specifies the prior mean outcome model size $m_{L}$. If not stated otherwise, we use $m_{L}=p/2$ as our default choice. For the treatment model we use a similar procedure, leading to uniform hyperpriors on the prior inclusion probabilities, as in Scott and Berger, (2010).
@@
A related question is what prior is induced on the number of valid and relevant instruments. This is described in supplementary Subsection S.3.2 and compared with the prior implied by IVBMA (with $l=1$).
@@ 4.1 Computational strategy
    Sample the covariance matrix $\bm{\Sigma}^{(s)}$ from its full conditional distribution.
@@
where $h$ is a proposal kernel. Throughout, models are proposed by randomly permuting the inclusion index of one covariate. This proposal is symmetric in the sense that $h(L\mid L^{\prime})/h(L^{\prime}\mid L)=1$, so the acceptance probability reduces to the (conditional) Bayes factor times the prior ratio (capped at 1). This proposal can lead to slow mixing in very high dimensions, where other proposals might be more attractive, but we leave this to future work.
@@ 4.2 Conditional Bayes factors and posteriors
Using the IW prior the conditional posterior for the covariance matrix is
$\bm{\Sigma}\mid\bm{\rho,\Lambda},L,M,\bm{y,X}\sim IW(\nu+n,\bm{I_{l+1}}+\bm{[\epsilon:H]}^{\intercal}\bm{[\epsilon:H]}).$ (8)
@@
For the Cholesky-based prior, we can use the conditional posteriors for the three covariance components,
$\bm{a_{yx}}^{\intercal}\mid\sigma_{y\mid x},\bm{\rho,\Lambda},L,M,\bm{y,X}$ $\sim N\left(\left(\bm{H}^{\intercal}\bm{H}+\frac{\sigma_{y\mid x}}{\omega_{a}}\bm{I_{l}}\right)^{-1}\bm{H}^{\intercal}\bm{\epsilon},\sigma_{y\mid x}\left(\bm{H}^{\intercal}\bm{H}+\frac{\sigma_{y\mid x}}{\omega_{a}}\bm{I_{l}}\right)^{-1}\right),$ (9)
$\sigma_{y\mid x}\mid\bm{a_{yx}},\bm{\rho,\Lambda},L,M,\bm{y,X}$ $\sim IG\left(\frac{n+\nu}{2},\frac{1}{2}\left(\left(\bm{\epsilon}-\bm{H}\bm{a_{yx}}^{\intercal}\right)^{\intercal}\left(\bm{\epsilon}-\bm{H}\bm{a_{yx}}^{\intercal}\right)+1\right)\right),$ (10)
$\bm{\Sigma_{xx}}\mid\bm{\rho,\Lambda},L,M,\bm{y,X}$ $\sim IW\left(n+\nu-1,\bm{H}^{\intercal}\bm{H}+\bm{I_{l}}\right).$ (11)
A draw of the covariance matrix $\Sigma$ is then implied by equation (4).
@@ 5 Model selection consistency of gIVBMA
  • under a hyperprior $p(g_{M})$, we have for $c\in\{l,2l,\dots,pl\}$
@@ 6 Simulation experiments
We evaluate the performance of gIVBMA in three different scenarios and compare it to naive BMA, IVBMA, and several classical methods. Throughout, we consider the median absolute error (MAE) and median bias of the point estimates of $\bm{\tau}$, (credible or confidence) interval coverage of $\bm{\tau}$, and the log predictive score (LPS) on a separate holdout dataset. For all Bayesian methods, we use the posterior mean as our point estimate for the MAE and bias calculations. While MAE, bias, and coverage relate specifically to the quality of the estimation of $\bm{\tau}$, LPS is a measure of predictive adequacy, focused on probabilistic predictions of the outcomes. In each scenario, we simulate $100$ independent datasets of size $n$, and for the LPS computation, we also generate holdout datasets of size $n/5$. Supplementary Section S.5.1 provides more details on the different measures used, while Section S.5.2 describes the different estimators and their implementation.
The predictive comparison is perhaps not entirely fair to some of the classical methods (in particular those that are two-stage in nature) as they reduce the bias in $\bm{\tau}$ but lack a correction term like gIVBMA. Consequently, when making predictions, these methods use an estimate of $\bm{\tau}$ instead of $\bm{\tau+\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$, which would be the appropriate coefficient for the conditional model $\bm{y}\mid\bm{X}$. This can lead to poor predictions when $\bm{\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$ becomes large. However, in practice, these methods might not be employed when prediction is the primary objective. Methods that do not correct for endogeneity (e.g., naive BMA) target the biased coefficient $\bm{\tau+\Sigma_{xx}^{-1}\Sigma_{yx}}^{\intercal}$, but provide good predictions as this is the right coefficient for predicting $\bm{y}$ conditional on $\bm{X}$. This discrepancy reflects the fact that predicting observed and counterfactual outcomes are different problems under endogeneity. Our probabilistic method allows us to provide endogeneity-corrected inference on $\bm{\tau}$ while maintaining good predictive power in the observational model.
@@ 6.1 Invalid instruments
To compare the performance of our different prior specifications, we evaluate several variants of our method using BRIC or hyper-$g/n$ priors on the regression coefficients, and either the full inverse-Wishart or the Cholesky-based covariance prior, the latter with different hyperparameters $\omega_{a}\in\{0.1,1,10\}$. We benchmark our approach against IVBMA (Karl and Lenkoski,, 2012), TSLS with all instruments, oracle TSLS, and the sisVIVE estimator (Kang et al.,, 2016). We also use a Bayesian IV method that includes all instruments in both models with horseshoe priors (Carvalho et al.,, 2009) on their regression coefficients. See supplementary Section S.5.2 for more details.
Table 1 reports the instrument selection performance of the different gIVBMA variants. The true number of valid and relevant instruments implied by the data-generating process is $N_{Z}=p-s$. For $n=50$, all variants distribute their posterior mass between $N_{Z}=0$ (exceeding the prior probability) and $0<N_{Z}<p-s$, placing only negligible mass on the true value (or above). When $s=3$, the mass is roughly evenly split, whereas for $s=6$, the models assign more weight to non-identified specifications. In contrast, for $n=500$, all variants allocate only negligible posterior mass to non-identified models. Overall, the hyper-$g/n$ specifications yield consistently better performance, while smaller values of $\omega_{a}$ seem to do better for $n=500$.
$n=50$
$s=3$ $s=6$
0 (0,ps)(0,p-s) psp-s (ps,p](p-s,p] 0 (0,ps)(0,p-s) psp-s (ps,p](p-s,p]
Prior probabilities 0.275 0.634 0.039 0.053 0.275 0.43 0.085 0.211
gIVBMA (BRIC, IW) 0.573 0.427 0.0 0.0 0.807 0.183 0.005 0.004
gIVBMA (h-$g/n$, IW) 0.488 0.508 0.003 0.0 0.732 0.235 0.017 0.016
gIVBMA (BRIC, $\omega_{a}=0.1$) 0.603 0.397 0.0 0.0 0.815 0.184 0.001 0.0
gIVBMA (h-$g/n$, $\omega_{a}=0.1$) 0.488 0.507 0.005 0.0 0.706 0.285 0.008 0.0
gIVBMA (BRIC, $\omega_{a}=1$) 0.595 0.405 0.0 0.0 0.779 0.195 0.007 0.019
gIVBMA (h-$g/n$, $\omega_{a}=1$) 0.51 0.487 0.003 0.0 0.705 0.224 0.017 0.054
gIVBMA (BRIC, $\omega_{a}=10$) 0.58 0.42 0.0 0.0 0.735 0.236 0.022 0.008
gIVBMA (h-$g/n$, $\omega_{a}=10$) 0.494 0.503 0.003 0.0 0.651 0.19 0.036 0.123
$n=500$
$s=3$ $s=6$
0 (0,ps)(0,p-s) psp-s (ps,p](p-s,p] 0 (0,ps)(0,p-s) psp-s (ps,p](p-s,p]
gIVBMA (BRIC, IW) 0.053 0.495 0.453 0.0 0.159 0.537 0.304 0.0
gIVBMA (h-$g/n$, IW) 0.05 0.417 0.533 0.0 0.158 0.468 0.355 0.019
gIVBMA (BRIC, $\omega_{a}=0.1$) 0.001 0.529 0.471 0.0 0.038 0.606 0.355 0.0
gIVBMA (h-$g/n$, $\omega_{a}=0.1$) 0.001 0.419 0.58 0.0 0.034 0.542 0.424 0.0
gIVBMA (BRIC, $\omega_{a}=1$) 0.05 0.498 0.452 0.0 0.205 0.507 0.288 0.0
gIVBMA (h-$g/n$, $\omega_{a}=1$) 0.065 0.397 0.538 0.0 0.168 0.446 0.352 0.034
gIVBMA (BRIC, $\omega_{a}=10$) 0.083 0.474 0.443 0.0 0.278 0.463 0.257 0.003
gIVBMA (h-$g/n$, $\omega_{a}=10$) 0.098 0.393 0.509 0.0 0.227 0.401 0.271 0.101
Table 1: Invalid instruments: Mean posterior probabilities of the number of valid and relevant instruments $N_{Z}$ for different variants of gIVBMA across $100$ simulated datasets with $n$ observations and $s$ invalid instruments. The true number of valid and relevant instruments is $N_{Z}=p-s$.
Table 2 shows that all gIVBMA variants predict well and lead to small bias and MAE. For small $n$, the BRIC specifications mostly have a higher MAE and median bias than the hyper-$g/n$ variants. The IVBMA method also predicts well and has low MAE and median bias, but overcovers for small $n$ and significantly undercovers for large $n$. TSLS results in a very large positive bias, very small coverage, and poor prediction. The oracle version of TSLS does better but is slightly worse than the gIVBMA methods for small $n$. All the methods above appear largely unaffected by the plurality rule. The latter does affect sisVIVE, however, which performs very poorly when $s=6$, as expected. The Bayesian horseshoe (BayesHS) method incurs substantial bias and does not cover well at all.
@@
BMA (hyper-g/n) 0.47 0.47 0.1 1.39 0.47 0.47 0.08 1.41
gIVBMA (BRIC, IW) 0.3 0.29 0.93 1.39 0.24 0.22 0.99 1.41
gIVBMA (hyper-$g/n$, IW) 0.14 0.06 0.9 1.4 0.17 0.01 0.99 1.41
gIVBMA (BRIC, $\omega_{a}=0.1$) 0.42 0.42 0.85 1.38 0.42 0.42 0.95 1.41
gIVBMA (hyper-$g/n$, $\omega_{a}=0.1$) 0.33 0.33 0.92 1.39 0.33 0.33 0.99 1.41
gIVBMA (BRIC, $\omega_{a}=1$) 0.24 0.23 0.9 1.39 0.26 0.2 0.97 1.41
gIVBMA (hyper-$g/n$, $\omega_{a}=1$) 0.18 0.01 0.91 1.4 0.24 0.09 0.95 1.41
gIVBMA (BRIC, $\omega_{a}=10$) 0.47 0.02 0.83 1.39 0.59 0.03 0.92 1.41
gIVBMA (hyper-$g/n$, $\omega_{a}=10$) 0.33 0.21 0.86 1.4 0.62 0.28 0.81 1.41
BayesHS 0.37 0.37 0.38 1.78 0.49 0.49 0.37 1.83
TSLS 1.25 1.25 0.18 2.17 2.43 2.43 0.09 2.41
O-TSLS 0.21 0.2 0.79 1.43 0.3 0.27 0.89 1.54
IVBMA 0.08 0.04 0.99 1.37 0.13 0.12 1.0 1.46
sisVIVE 0.6 0.6 - 1.64 2.2 2.2 - 2.37
@@
BMA (hyper-g/n) 0.46 0.46 0.0 1.3 0.48 0.48 0.0 1.28
gIVBMA (BRIC, IW) 0.1 0.06 0.98 1.29 0.21 0.2 0.97 1.28
gIVBMA (hyper-$g/n$, IW) 0.08 0.06 0.94 1.29 0.21 0.18 0.94 1.28
gIVBMA (BRIC, $\omega_{a}=0.1$) 0.1 0.09 0.93 1.29 0.16 0.16 0.89 1.28
gIVBMA (hyper-$g/n$, $\omega_{a}=0.1$) 0.09 0.08 0.92 1.29 0.17 0.17 0.89 1.28
gIVBMA (BRIC, $\omega_{a}=1$) 0.08 0.06 0.93 1.29 0.2 0.18 0.96 1.28
gIVBMA (hyper-$g/n$, $\omega_{a}=1$) 0.1 0.07 0.95 1.29 0.19 0.18 0.93 1.28
gIVBMA (BRIC, $\omega_{a}=10$) 0.1 0.07 0.97 1.29 0.32 0.31 0.89 1.28
gIVBMA (hyper-$g/n$, $\omega_{a}=10$) 0.11 0.08 0.93 1.29 0.52 0.52 0.7 1.28
BayesHS 0.35 0.35 0.04 1.34 0.51 0.51 0.02 1.3
TSLS 1.82 1.82 0.0 2.2 3.63 3.63 0.0 2.7
O-TSLS 0.07 0.04 0.92 1.42 0.08 0.05 0.96 1.4
IVBMA 0.09 0.08 0.78 1.28 0.09 0.06 0.74 1.3
sisVIVE 0.21 0.21 - 1.61 4.36 4.36 - 3.07
Table 2: Invalid instruments: MAE, median bias, coverage, and mean LPS on 100 simulated datasets with $s$ invalid instruments. The best values in each column are printed in bold. The sisVIVE estimator does not provide any uncertainty quantification, so we do not report any coverage results.
In the subsequent simulations, we shall focus on the hyper-$g/n$ prior on the regression coefficients, using either the full inverse-Wishart covariance prior or the Cholesky-based prior with $\omega_{a}=0.1$, as these performed best in this experiment. The Cholesky prior with $\omega_{a}=1$ also performed well, but behaved similarly to the full inverse-Wishart prior. This is not surprising, since $\sigma_{y\mid x}$ is roughly of the order 1 in the IW prior (see Supplementary Figure S.1).
@@ 6.2 Multiple endogenous variables with correlated instruments
We compare gIVBMA against naive BMA, IVBMA, OLS, TSLS (using all instruments and an oracle version), and MATSLS. The latter refers to the model-averaged TSLS estimator with unrestricted weights by Kuersteiner and Okui, (2010) and is included as it is the only feasible classical method that is competitive with oracle TSLS in all scenarios of Subsection 6.3. IVBMA cannot directly model the Beta-distributed endogenous variable and instead relies on a Gaussian approximation.
Supplementary Table S.1 presents the full estimation error, coverage, and LPS results. Both retained gIVBMA variants perform best in terms of MAE and median bias, for both sample sizes. BMA is severely biased as expected, but coverage for the second endogenous variable is adequate (in contrast with the coverage for $\bm{X}_{1}$). This is a consequence of the non-linear transformation from the latent Gaussian to the Beta variable, diluting the endogeneity. TSLS, oracle TSLS, and MATSLS all perform similarly well. IVBMA results in a substantial bias for both sample sizes, yet it predicts and covers relatively well.
@@
Figure 1 presents the distributions of the posterior probability of the true treatment model and the mean treatment model size. Unlike gIVBMA, IVBMA allows for separate treatment models for the two endogenous variables (see Subsection 2.2), and we report the results separately for $X_{1}$ and $X_{2}$. Both gIVBMA variants put substantial posterior mass on the true model, and their mean model size is accordingly very close to that of the true model (which is $5$). For $n=50$, IVBMA tends to select too many variables for $X_{1}$ and too few for $X_{2}$. The performance of IVBMA is better in the $n=500$ setting, but it still cannot compete with gIVBMA. Supplementary Table S.2 presents median posterior inclusion probabilities (PIP) for all instruments. The selection performance of gIVBMA is clearly superior for both sample sizes.